A jug is at 15oC and is filled with water at an initial temperature of 85oC. The mass of water is 2.5kg. SHC is 4200. If the final temperature of the water is 34oC:

1. what is the final temperature of the jug?
2. if the jug has a SHC of 400J/kg/oC, what is its mass?
1
Expert's answer
2020-12-06T17:16:48-0500

"t_1=15°C, ~t_2=85°C, ~m_2=2.5 kg, ~c_1=400 \frac{J}{kg\cdot °C},~c_2=4200 \frac{J}{kg\cdot °C},~t=34°C."

"m_1-?"


a)

Jug's temperature is equal to water's temperature by thermodynamic equilibrium.

b)

"Q_1=-Q_2."

"Q_1=c_1m_1(t-t_1),"

"Q_2=c_2m_2(t-t_2)."

"m_1=\frac{c_2m_2(t_2-t)}{c_1(t-t_1)},"

"m_1=\frac{4200\cdot2.5\cdot(85-34)}{400\cdot(34-15)}=70.5~kg."